OpenStudy (zale101):

Vector problem: I’m trying to determine if the mentioned vertices are right triangle.

OpenStudy (zale101):

A(4,5-3) B(3,2,-1) C(7,0,-2) AB=(3-4)i+(2-5)j+(-1+3)k =-i-3j+2k=<-1,-3,2> BC=(7-3)i+(0-2)j+(-2+1)k =4i-2j-k=<4,-2,-1> AC=(7-4)i+(0-5)j+(-2+3)k =3i-5j+k=<3,-5,1> AB*BC=-1*4+-3*-2+2*-1=1 AB*AC=-1*3+-3*-5+2*1=14 BC*AC=4*3+-2*-5+-1*1=21 (pardon my missing vector arrows, and hats on i,j,k)

OpenStudy (zale101):

Somehow, by using the dot product, i couldn't get zero in either AB*BC, AB*AC, and BC*A

OpenStudy (zale101):

There's an error with my arithmetic i suppose.

OpenStudy (dan815):

have u learnt how to take the dot product

jimthompson5910 (jim_thompson5910):

you did make an error when you did the dot product

OpenStudy (dan815):

either that or see the length are in the form a^2+b^2=c^2

OpenStudy (zale101):

Oh okay. Thanks! This semester is like my first time learning the doc product, can't remember it from precal. I'll go over what i did. Thanks again @dan815 and @jim_thompson5910 .

jimthompson5910 (jim_thompson5910):

you have the right steps, just a silly arithmetic mistake

OpenStudy (zale101):

Lol Alright, i feel good now. I was afraid that i didn't know what i'm trying to solve for.

OpenStudy (dan815):

the dot product and the cow law are very closely related

OpenStudy (zale101):

AC*BC=-1*4+-3*-2+2*-1=0

jimthompson5910 (jim_thompson5910):

I think you mean AB*BC=-1*4+-3*-2+2*-1=0 right?

OpenStudy (zale101):

Yes that AB and BC

OpenStudy (zale101):

Sorry xD

jimthompson5910 (jim_thompson5910):

that's ok. I'm getting the same result. So AB is orthogonal (ie perpendicular) to BC